# infocolor=gray

### v2

@@ Abstract
Kolmogorov–Arnold Networks (KANs) have shown strong potential for efficiently approximating complex nonlinear functions. However, the original KAN formulation relies on B-spline basis functions, which incur substantial computational overhead due to De Boor’s algorithm. To address this limitation, recent work has explored alternative basis functions such as radial basis functions (RBFs) that can improve computational efficiency and flexibility. Yet, standard RBF-KANs often sacrifice accuracy relative to the original KAN design. In this work, we propose Free-RBF-KAN, a RBF-based KAN architecture that incorporates adaptive learning grids and trainable smoothness to close this performance gap. Our method employs freely learnable RBF shapes that dynamically align grid representations with activation patterns, enabling expressive and adaptive function approximation. Additionally, we treat smoothness as a kernel parameter optimized jointly with network weights, without increasing computational complexity. We provide a general universality proof for RBF-KANs, which encompasses our Free-RBF-KAN formulation. Through a broad set of experiments, including multiscale function approximation, physics-informed machine learning, and PDE solution operator learning, Free-RBF-KAN achieves accuracy comparable to the original B-spline-based KAN while delivering faster training and inference. These results highlight Free-RBF-KAN as a compelling balance between computational efficiency and adaptive resolution, particularly for high-dimensional structured modeling tasks.
@@ 1 Introduction
The Kolmogorov–Arnold Network (KAN) [Liu et al., 2024] is a neural architecture grounded in the Kolmogorov–Arnold representation theorem [Kolmogorov, 1956], which states that any multivariate continuous function can be expressed as a superposition of univariate continuous functions and addition. The original KAN leverages B-splines to model these univariate components due to their strong approximation capabilities. KAN has been shown to exhibit no spectral bias [Wang et al., 2024a], making it advantageous for learning functions with high-frequency components. However, in practice, computing B-spline bases using De Boor’s iteration and the required rescaling of the B-spline domain during training introduces significant computational overhead. To mitigate this issue, recent work has explored alternative basis functions and approaches. Li [2024] proposed FastKAN, which replaces B-splines with Gaussian radial basis functions (RBFs) and uses layer normalization to avoid domain rescaling, offering a simpler and faster implementation while retaining competitive accuracy. Extending this approach, FasterKAN [Delis, 2024] introduces a reflectional switch activation function (RSWAF) that approximates Gaussian RBFs without costly exponential evaluations, leading to further reductions in computational cost. Adaptive meshing via FreeKnots-KAN, as proposed in Zheng et al. [2025] and Actor et al. [2025], improved both the accuracy and stability by learning the placement of B-spline knots.
All of the aforementioned modifications to the basic KAN architecture can be viewed as extensions of radial basis function (RBF) networks [Orr and others, 1996], which replace the traditional sigmoidal nonlinearities in neural networks (NNs) with RBFs. RBF networks are universal function approximators and have been shown to offer an accurate and interpretable alternative to standard NNs [Montazer et al., 2018]. However, they do not scale well to high-dimensional problems: the RBFs themselves become unwieldy high-dimensional functions, and the number of grid centers grows exponentially with the input dimension, reflecting the classic curse of dimensionality. In contrast, RBF-KAN leverages univariate RBF basis functions within the superpositional structure prescribed by the Kolmogorov–Arnold theorem, enabling scalable function approximation even in high-dimensional settings.
@@
In this paper, we propose Free-RBF-KAN, which adopts a hierarchical, multichannel structure rooted in the Kolmogorov–Arnold decomposition. The Free-RBF-KAN architecture integrates adaptive learning grids (i.e., free knots) with tunable RBF shape parameters. This formulation constrains the grid to a fixed domain while learning a mesh aligned with activation patterns. The resulting geometric refinement balances computational efficiency and expressiveness, making it well-suited for modeling structured, high-dimensional functions with varying local complexity. Each RBF kernel in Free-RBF-KAN acts as a univariate component, and their superposition constructs the multivariate function, aligning with the Kolmogorov–Arnold theorem. To enhance performance, Free-RBF-KAN uses trainable grids and a shape mechanism similar to Ramabathiran and Ramachandran [2021], Wettschereck and Dietterich [1991], Mojarrad et al. [2023]. Our method permits dynamic repositioning of both the grid points and the smoothness during training. The adaptive mesh has been integrated into KAN by Actor et al. [2025] and Zheng et al. [2025], and enhances resolution and adaptability by decoupling the mesh from fixed uniform structures. Furthermore, Actor et al. [2025] provides a theoretical justification for KAN that shows that spline-based KANs can serve as preconditioning for multichannel MLPs, which yields improved optimization landscapes and faster convergence.
@@
As a theoretical contribution of this work, we formally extend the universal approximation theorem for RBF networks to the RBF-KAN architecture. To the best of our knowledge, this is the first universality proof established for the RBF-KAN family. In contrast to KANs that rely on B-spline bases, RBF-KAN exhibits stronger function-approximation properties: its approximation error bound is independent of the target function, and it does not require the target function to admit any specific or predesigned decomposition. This inherent universality demonstrates that RBF-KAN is not merely a strategy for reducing computational overhead, but a fundamentally powerful and flexible framework for function approximation. Furthermore, we evaluate the regression performance of Free-RBF-KAN on high-dimensional dataset and analyze its Neural Tangent Kernel (NTK). Our NTK study confirms that Free-RBF-KAN, like the original KAN, does not exhibit spectral bias.
@@
We further explore the application of Free-RBF-KAN to physics-informed machine learning and operator learning tasks. Using a physics-informed loss in the spirit of Physics-Informed Neural Networks (PINNs) [Raissi et al., 2019], we train Free-RBF-KAN to solve a 1D heat conduction PDE and a 2D Helmholtz PDE. Notably, the PINN baseline fails to converge on the heat conduction problem, and on the Helmholtz benchmark it produces substantially larger errors than physics-informed Free-RBF-KAN, even though the PINN uses two orders of magnitude more parameters. We also evaluate Free-RBF-KAN as the trunk network within a DeepONet [Lu et al., 2019] for learning the solution operator of a reaction–diffusion PDE. In contrast to DeepOKAN [Abueidda et al., 2025], we retain a standard MLP for the branch network to maximize performance. Free-RBF-KAN achieves lower approximation error than the standard DeepONet while requiring fewer parameters. Across all experiments, Free-RBF-KAN consistently outperforms both RBF-KAN and the original KAN.
@@ 1.1 Main Contributions
The central contribution of this work is the development of Free-RBF-KAN, a novel and highly efficient Kolmogorov-Arnold Network architecture utilizing a flexible RBF formulation. Our work provides the following specific advancements:
@@ 2 Related Work
A wide range of works have extended the Kolmogorov–Arnold Network (KAN) and Radial Basis Function (RBF) frameworks in both theoretical and application-driven directions. Physics-Informed RBF networks using Gaussian kernels have demonstrated superior performance on high-frequency PDEs [Bai et al., 2023], while Zeng et al. [2024] proposed RBF-PINN as an alternative to Fourier embeddings. Within the KAN framework, SS et al. [2024] introduced Chebyshev basis functions, and Wang et al. [2024b] employed third-order B-splines in KAN to encode physical laws in strong, energy, and inverse forms. Hybrid models such as BSRBF-KAN [Ta, 2024] combine B-splines and RBFs to harness the benefits of both smooth local representation and adaptive flexibility. Krisnawan et al. [2025] combines the architectures of KAN and RBF network for accurate indoor localization using RSSI-based fingerprinting. Bai et al. [2023] demonstrated that physics-informed RBF networks can outperform traditional PINNs, though these models typically use a single-layer RBF architecture. Similarly, Ta [2024] combined B-spline and RBF bases to enhance training, while Shukla et al. [2024] and Abueidda et al. [2025] introduced DeepOKAN, which uses RBF-KAN for operator learning. Li [2024] demonstrates that B-Spline can be approximated by radial basis function with Gaussian Kernel. A thorough comparison of PINN and KAN-based models is provided by Shukla et al. [2024]. On the theoretical side, classical results on RBF networks include their universal approximation capabilities [Park and Sandberg, 1993, Ismayilova and Ismayilov, 2024], convergence and consistency properties [Xu et al., 1994], and optimal approximation results [Girosi and Poggio, 1990]. RBF networks have also been used for solving multiscale PDEs [Wang et al., 2023b]. Earlier work by Wettschereck and Dietterich [1991] focused on learning RBF centers to improve performance. Connections between RBFs and kernel machines were explored by Que and Belkin [2016], while Chen [2024] extended KAN to a Gaussian process formulation that provides error estimates. Theoretical refinements of Kolmogorov’s representation theorem include improvements by Fridman [1967], a constructive proof by Braun and Griebel [2009], and an extension by Kurkova [1991] showing that the representation can be realized through affine and sigmoidal functions—permitting the approximation of discontinuous but bounded functions. This theory underpins recent developments such as the KKAN model proposed in Toscano et al. [2025], which builds on the Kurkova–Kolmogorov–Arnold representation. A broader perspective on KAN developments is provided by Somvanshi et al. [2024], which highlights applications across scientific computing, time-series forecasting, and graph learning.
@@ 3.1 Kolmogorov-Arnold Theorem
The Kolmogorov-Arnold Network (KAN) [Liu et al., 2024] is inspired by the Kolmogorov-Arnold Theorem [Kolmogorov, 1956]. The theorem states that for any integer $n\geq 2$ there exist continuous univariate real functions $\phi^{(p,q)}(x):[0,1]\to\mathbb{R}$ and $\phi^{(q)}:\mathbb{R}\to\mathbb{R}$ that can represent continuous multivariate real function $f(x_{1},\dots,x_{d}):[0,1]^{d}\to\mathbb{R}$ as
$f(x_{1},\dots,x_{d})=\sum_{q=1}^{2d+1}\phi^{(q)}\left(\sum_{p=1}^{d}\phi^{(p,q)}(x_{p})\right)\,.$ (1)
This result provides a universal representation for high-dimensional functions using only sums and compositions of univariate functions. However, the decomposition (1) may involve non-smooth inner functions $\phi^{(p,q)}$, and is hard for exact representation as shown in Girosi and Poggio [1989].
3.2 Radial Basis Functions
Radial basis functions (RBFs) are real-valued functions whose output depends solely on the distance from a central point. RBF networks are widely used in interpolation, approximation theory, and machine learning, and classical results (e.g., [Park and Sandberg, 1991]) show that they are universal approximators of continuous functions on compact domains. In the one-dimensional setting, a single-layer RBF network is takes the form
@@ 3.2 Radial Basis Functions
where $G\in\mathbb{N^{+}}$ is the number of nodes, $x\in\mathbb{R}$ is the input, the kernel $K:\mathbb{R}\rightarrow\mathbb{R}^{+}$ depends only on $|x-c_{m}|$. Each term of (2) is parameterized by a weight $\omega_{m}$, a centroid $c_{m}$, and a smoothness factor $\sigma_{m}>0$. In this work, we focus on one-dimensional RBFs as fundamental building blocks for approximating multivariate functions through the superpositional structure provided by the Kolmogorov–Arnold Theorem.
@@ 4 Free RBF-KAN
The original KAN [Liu et al., 2024] is loosely inspired by the Kolmogorov–Arnold theorem. Assuming the target function is sufficiently smooth, KAN employs B-spline basis functions for the functions $\phi^{(p,q)}$ in (1). Its expressivity derives from a multilayer architecture—an extension not present in the original formulation of Kolmogorov [1956], but one that substantially improves practical performance. RBF-KAN preserves the general structure of [Liu et al., 2024], replacing the B-spline bases with one-dimensional radial basis function (RBF) kernels. Let $n_{l}$ denote the number of nodes in the $l$-th layer and $\textbf{x}^{(l)}$ the inputs to layer $l$. The multilayer RBF-KAN then satisfies the following recursive relation:
@@
for which $K\in C^{\infty}(\mathbb{R})$. Another widely used option is the Matérn kernel with smoothness parameter, such as $\nu=5/2$ which satisfies $K\in C^{3}(\mathbb{R})$. The smoothness of the chosen kernel directly determines the smoothness of the resulting RBF-KAN output. In this study, we investigate trainable centroids and smoothness parameters to provide additional flexibility and improve representation quality. Throughout the paper, RBF-KAN refers to networks with fixed centroids and fixed smoothness parameters, and Free-RBF-KAN refers to models in which both centroids and smoothness factors are learnable.
@@ 4.1 Adaptive Meshing
The free-knots method for KAN has been explored in Zheng et al. [2025] and Actor et al. [2025]. In this work, we extend the underlying idea to the RBF-KAN setting. Unlike B-splines, which require maintaining a strictly ordered sequence of grid points, RBFs introduce additional flexibility of allowing any order of centroids, as each kernel is evaluated independently and does not rely on a prescribed ordering of centroids. This allows us to develop a moving grid method in the RBF framework that can adaptively remesh the activation functions during training. It is worth noting that the free-knot adaptation for B-splines is considerably more complicated and the smoothness of B-spline bases cannot be treated as a trainable parameter. In contrast, RBFs enable both the centroid locations and the associated smoothness parameters to be learned directly and efficiently, yielding a more expressive and computationally efficient approach.
To obtain an efficient representation, we constrain the grids to lie within a prescribed domain during training. This can be achieved by reparameterization using a bounded monotonic activation function $\rho:\mathbb{R}\to[a,b]$ with $a<b$ and $\rho\in C^{\infty}$. Given a grid domain $\Omega_{g}\in(x_{l},x_{r})$ with $x_{l}<x_{r}$, a free parameter $\tilde{c}\in\mathbb{R}$ is mapped to a valid centroid location $c\in(x_{l},x_{r})$ via
$c=x_{l}+\frac{x_{r}-x_{l}}{b-a}\left(\rho(\tilde{c})-a\right)$ (6)
In this work, without specification, we set a=1,b=1a=-1,b=1, and choose $\rho$ to be the tanh function, ensuring that centroid $c$ remains with in $\Omega_{g}$. This reparameterization is smooth and compatible with gradient-based optimization. For initialization, the grid points are placed uniformly within the domain.
@@ 4.2 Adaptive Smoothness
The smoothness of B-splines (determined by their polynomial orders) is fixed and must be predetermined prior to the training. In contrast, the smoothness of RBFs can be specified either by the order parameter $\nu$ in the case of the Matèrn kernel, or more generally by the smoothness factor $\sigma$. To ensure that $\sigma$ remains positive during gradient-based optimization, we introduce an unconstrained parameter $\tilde{\sigma}\in\mathbb{R}$ and define $\sigma$ via the mapping
@@
The combination of adaptive meshing and adaptive smoothness can significantly enhance the express power of Free-RBF-KAN. As demonstrated in Zheng et al. [2025], allowing the RBF centroids to move within an extended range improves gradient smoothness and helps training stability. Although the introduction of additional trainable parameters may moderately increase the training time, Free-RBF-KAN remains substantially faster than B-spline KAN. Moreover, the inference cost of Free-RBF-KAN is identical to that of standard RBF-KAN once the centroids and smoothness parameters have been trained and fixed.

### v3

# Free-RBF-KAN: Kolmogorov–Arnold Networks with Adaptive Radial Basis Functions for Efficient Function Learning

@@ Abstract
Kolmogorov–Arnold Networks (KANs) offer a promising framework for approximating complex nonlinear functions, yet the original B-spline formulation suffers from significant computational overhead due to De Boor’s algorithm. While recent RBF-based variants improve efficiency, they often sacrifice the approximation accuracy inherent in the original spline-based design. To bridge this gap, we propose Free-RBF-KAN, an architecture that integrates adaptive learning grids and trainable smoothness parameters to enable expressive, high-resolution function approximation. Our method utilizes learnable RBF shapes that dynamically align with activation patterns, and we provide the first formal universal approximation proof for the RBF-KAN family. Empirical evaluations across multiscale regression, physics-informed PDEs, and operator learning demonstrate that Free-RBF-KAN can achieve accuracy comparable to its B-spline counterparts while delivering significantly faster training and inference. These results establish Free-RBF-KAN as an efficient and adaptive alternative for high-dimensional structured modeling tasks.
@@ 1 Introduction
The Kolmogorov–Arnold Network (KAN) [Liu et al., 2024] is a neural architecture grounded in the Kolmogorov–Arnold representation theorem [Kolmogorov, 1956], which states that any multivariate continuous function can be expressed as a superposition of univariate continuous functions and addition. The original KAN leverages B-splines to model these univariate components due to their strong approximation capabilities and lack of spectral bias [Wang et al., 2024a]. However, computing B-spline bases using De Boor’s iteration and the necessary domain rescaling during training introduces significant computational overhead. To mitigate this, recent works, e.g., FastKAN [Li, 2024] and FasterKAN [Delis, 2024] have explored alternative basis functions, such as Gaussian radial basis functions (RBFs) and reflectional switch activation functions (RSWAF), to reduce costs while retaining accuracy. Further improvements in stability and accuracy have been achieved through adaptive meshing in FreeKnots-KAN [Zheng et al., 2025], which learns the optimal knot placement. These modifications can be viewed as extensions of RBF networks [Orr et al., 1996], which replace traditional sigmoidal nonlinearities with RBFs. While RBF networks are universal function approximators, they typically struggle with the curse of dimensionality, as the number of grid centers required to achieve uniform approximation accuracy grows exponentially with input dimension. In contrast, RBF-KAN leverages univariate RBF basis functions within the superpositional structure of the Kolmogorov–Arnold theorem, enabling scalable function approximation in high-dimensional settings.
@@
In this paper, we propose Free-RBF-KAN, which adopts a hierarchical, multichannel structure integrating adaptive learning grids (free knots) with tunable RBF shape parameters. This architecture constrains the grid to a fixed domain while learning a mesh aligned with activation patterns, balancing computational efficiency and expressiveness. Each RBF kernel acts as a univariate component whose superposition constructs the multivariate function. By permitting dynamic repositioning of grid points and smoothness during training, Free-RBF-KAN decouples the mesh from fixed uniform structures. Furthermore, our approach aligns with findings that spline-based KANs can serve as preconditioning for multichannel multilayer perceptrons (MLPs), leading to improved optimization landscapes and faster convergence [Actor et al., 2025].
@@
As a theoretical contribution, we formally extend the universal approximation theorem for RBF networks to the RBF-KAN architecture. To our knowledge, this is the first universality proof for the RBF-KAN family. Unlike B-spline KANs, RBF-KAN exhibits an approximation error bound independent of the target function and does not require a predesigned decomposition. This inherent universality demonstrates that RBF-KAN is a fundamentally powerful framework rather than just a strategy for reducing overhead. We also analyze the Neural Tangent Kernel (NTK) of Free-RBF-KAN, confirming that it does not exhibit spectral bias.
@@
Finally, we evaluate Free-RBF-KAN on high-dimensional regression, physics-informed neural networks (PINNs), and operator learning. Using a physics-informed loss [Raissi et al., 2019], we solve heat conduction and Helmholtz PDEs. Notably, the PINN baseline fails to converge on the heat equation and produces larger errors on the Helmholtz benchmark, despite using two orders of magnitude more parameters. We also evaluate Free-RBF-KAN as the trunk network within a DeepONet [Lu et al., 2019] for learning solution operators of reaction–diffusion PDEs. Free-RBF-KAN achieves lower approximation error than standard DeepONets and DeepOKANs while requiring fewer parameters. Across all experiments, Free-RBF-KAN consistently outperforms both RBF-KAN and the original KAN.
@@ 1.1 Main Contributions
The central contribution of this work is the development of Free-RBF-KAN, a novel and highly efficient KAN architecture utilizing a flexible RBF formulation. Our work provides the following specific advancements:
@@ 2 Related Work
A wide range of studies have explored both theoretical developments and applications of KAN and RBF frameworks. Physics-informed RBF networks using Gaussian kernels have demonstrated superior performance on high-frequency PDEs [Bai et al., 2023], while Zeng et al. [2024] proposed RBF-PINN as an alternative to Fourier embeddings. SS et al. [2024] introduced Chebyshev basis functions within KAN, while Wang et al. [2024b] employed cubic B-splines to encode physical laws in strong, energy, and inverse forms. Hybrid models such as BSRBF-KAN [Ta, 2024] combine B-splines and RBFs to harness the benefits of both smooth local representation and adaptive flexibility. Krisnawan et al. [2025] combines KAN and RBF networks for accurate indoor localization using RSSI-based fingerprinting. Bai et al. [2023] demonstrated that physics-informed RBF networks can outperform traditional PINNs, using a single RBF layer. Similarly, Ta [2024] combined B-spline and RBF bases to enhance training, while Shukla et al. [2024] and Abueidda et al. [2025] introduced DeepOKAN, which employs RBF-KAN for operator learning. Li [2024] demonstrates that B-Spline can be approximated by RBFs with Gaussian kernels. A thorough comparison of PINN and KAN-based models is provided by Shukla et al. [2024].
On the theoretical side, classical results on RBF networks include their universal approximation capabilities [Park and Sandberg, 1993, Ismayilova and Ismayilov, 2024], convergence and consistency properties [Xu et al., 1994], and optimal approximation results [Girosi and Poggio, 1990]. RBF networks have also been used for solving multiscale PDEs [Wang et al., 2023b]. Earlier work by Wettschereck and Dietterich [1991] focused on learning RBF centers to improve performance. Connections between RBFs and kernel machines were explored by Que and Belkin [2016], while Chen [2024] extended KAN to a Gaussian process formulation that provides error estimates. Theoretical refinements of Kolmogorov’s representation theorem include the improvement by Fridman [1967], a constructive proof by Braun and Griebel [2009], and an extension by Kurkova [1991] showing that the representation can be realized through affine and sigmoidal functions–permitting the approximation of discontinuous but bounded functions. This theory underpins recent developments such as the KKAN model proposed in Toscano et al. [2025], which builds on the Kurkova–Kolmogorov–Arnold representation. A broader perspective on KAN developments is provided by Somvanshi et al. [2024], highlighting applications across scientific computing, time-series forecasting, and graph learning.
@@ 3.1 Kolmogorov-Arnold Theorem
KAN [Liu et al., 2024] is inspired by the Kolmogorov-Arnold representation theorem [Kolmogorov, 1956]. This theorem states that every multivariate continuous function $f\colon[0,1]^{n}\to\mathbb{R}$ can be represented as a superposition of continuous univariate functions. However, the superposition may involve non-smooth inner functions and is hard for exact representation Girosi and Poggio [1989].
3.2 Radial Basis Functions (RBFs)
RBFs are real-valued functions whose output depends solely on the distance from a central point. RBF networks are widely used in interpolation, approximation theory, and machine learning. Classical results [Park and Sandberg, 1991] show that RBFs are universal approximators of continuous functions on compact domains. In the one-dimensional setting, a single-layer RBF network takes the form
@@ 3.2 Radial Basis Functions (RBFs)
where $G\in\mathbb{N^{+}}$ is the number of nodes and kernel $K:\mathbb{R}\rightarrow\mathbb{R}^{+}$ depends only on $|x-c_{m}|$. Each term of (1) is parameterized by a weight $\omega_{m}$, a centroid $c_{m}$, and a smoothness factor $\sigma_{m}>0$. In this work, we focus on one-dimensional RBFs as the fundamental building blocks for approximating multivariate functions through the superpositional structure provided by the Kolmogorov–Arnold Theorem.
@@ 4 Free RBF-KAN
Assuming that the target function is sufficiently smooth, KAN employs B-spline basis functions for functions $\phi^{(p,q)}$ in Lemma 4.1. Its expressivity derives from a multilayer architecture, an extension not present in the original formulation of Kolmogorov [1956] but one that substantially improves practical performance. RBF-KAN preserves the structure of Liu et al. [2024], while replacing the B-splines with one-dimensional RBFs. Let $n_{l}$ denote the number of nodes in the $l$-th layer and $\textbf{x}^{(l)}$ the inputs to the layer. A multilayer RBF-KAN satisfies the following recursive relation:
@@
for which $K\in C^{\infty}(\mathbb{R})$. Another widely used option is the Matérn kernel with smoothness parameter, such as $\nu=5/2$ which satisfies $K\in C^{3}(\mathbb{R})$. The smoothness of the chosen kernel directly determines the smoothness of the resulting RBF-KAN. In this work, we investigate trainable centroids and smoothness parameters for additional flexibility and better representation quality. Throughout the paper, RBF-KAN refers to networks with fixed centroids and smoothness parameters, whereas Free-RBF-KAN denotes models where both centroids and smoothness factors are learnable.
@@ 4.1 Adaptive Meshing
Free-knots methods for KAN have been explored in Zheng et al. [2025] and Actor et al. [2025]. In this work, we extend the underlying idea to the RBF-KAN setting. Unlike B-splines which require maintaining a strictly ordered sequence of grid points, RBFs introduce additional flexibility of allowing any order of centroids, as each kernel is evaluated independently and does not rely on a prescribed ordering of centroids. This allows us to develop a moving grid method in the RBF framework that can adaptively remesh the activation functions during training. It is worth noting that the free-knot adaptation for B-splines is considerably more complicated and the smoothness of the B-splines, determined by their polynomial orders, is fixed and cannot be treated as a trainable parameter. In contrast, RBFs enable both the centroid locations and the associated smoothness parameters to be learned directly and efficiently, yielding a more expressive and computationally efficient approach.
To obtain an efficient representation, we constrain the grids to lie within a prescribed domain during training. This can be achieved by reparameterization using a bounded monotonic activation function $\rho:\mathbb{R}\to[a,b]$ with $a<b$ and $\rho\in C^{\infty}$. Given a grid domain $\Omega_{g}\in(x_{l},x_{r})$ with $x_{l}<x_{r}$, a free parameter $\tilde{c}\in\mathbb{R}$ is mapped to a valid centroid location $c\in(x_{l},x_{r})$ via $c=x_{l}+\frac{x_{r}-x_{l}}{b-a}\left(\rho(\tilde{c})-a\right)$. Unless otherwise specified, we set a=1,b=1a=-1,b=1, and choose $\rho$ to be the tanh function, ensuring that the centroid $c$ remains within $\Omega_{g}$. This reparameterization is smooth and compatible with gradient-based optimization. For initialization, the grid points are placed uniformly within the domain.